Find the angle between the two vectors
A= 2i +3j +k
B = I - 6j +k
1
Expert's answer
2019-02-06T15:25:09-0500

We can find the angle between the two vectors from the formula:

"cos\alpha = \frac{\overrightarrow{A} \cdot \overrightarrow{B}}{\lvert \overrightarrow{A} \rvert \cdot \lvert \overrightarrow{B} \rvert},"

here,

"\overrightarrow{A} \cdot \overrightarrow{B}"

is the dot product of two vectors

"\overrightarrow{A}"

and

"\overrightarrow{B}"

;

"\lvert \overrightarrow{A} \rvert"

,

"\lvert \overrightarrow{B} \rvert"

are the magnitudes of the vectors

"\overrightarrow{A}"

and

"\overrightarrow{B}"

, respectively.

Let's first find the dot product of two vectors

"\overrightarrow{A}"

and

"\overrightarrow{B}"

:

"\overrightarrow{A} \cdot \overrightarrow{B} = A_x \cdot B_x + A_y \cdot B_y + A_z \cdot B_z = 2 \cdot 1 + 3 \cdot (-6) + 1 \cdot 1 = -15."

Then, let's find the magnitudes of the vectors

"\overrightarrow{A}"

and

"\overrightarrow{B}"

:

"\lvert \overrightarrow{A} \rvert =\sqrt{A_x^2 + A_y^2 + A_z^2} = \sqrt{2^2 + 3^2 + 1^2} = \sqrt{14},""\lvert \overrightarrow{B} \rvert =\sqrt{B_x^2 + B_y^2 + B_z^2} = \sqrt{1^2 + (-6)^2 + 1^2} = \sqrt{38}."

Finally, we can calculate the angle between two vectors

"\overrightarrow{A}"

and

"\overrightarrow{B}"

:

"\alpha = arccos(\frac{\overrightarrow{A} \cdot \overrightarrow{B}}{\lvert \overrightarrow{A} \rvert \cdot \lvert \overrightarrow{B} \rvert}),""\alpha = arccos(\frac{-15}{\sqrt{14} \cdot \sqrt{38}}) = 130^ \circ."

Answer:

"\alpha = 130^ \circ"

.